craig is 5 years older than bruce. the product of their ages is 1184. how old are they? i know their ages are 32 and 37 but how do i set it up?
Answer by Maths68(1474)

Let
c=Current age of Craig
b=Current age of Bruce
If we add 5 years to the Bruce age then both will be of same age
c=b+5…….(1)
the product of their ages is 1184
c*b=1184
Divide above equation by ‘c’ both sides
cb/c=1184/c
b=1184/c……….(2)
Put the value of b in (1) we have
c=(1184/c)+5
Multiply by ‘c’ both sides
c^2=1184+5c
Now solve for ‘c’
c^2-5c-1184=0
c^2-37c+32c-1184=0
c(c-37)+32(c-37)=0
(c-37)(c+32)=0
c-37=0 or c+32=O
c=37 or c=-32 (which is impossible - Craig will not born after 32 years from now)
Therefore C=37 so Craig is 37 years old
Put the value of ‘c’ in equation (1) we have
c=b+5
37=b+5
37-5=b
b=32
So Bruce is 32 years old.